OpenStudy (anonymous):

There are three geometric means between 10 and 2,560. What is the value of tn? Select one: a. 10 b. 150 c. 1150 d. 2560

OpenStudy (anonymous):

@zepdrix

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 10 \times r^4=2,560\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

Solve for the common ratio, \(\normalsize\color{blue}{ 10 \times r^4=2,560\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ r^4=256\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ r^4=4^4\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ r=4\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

The knowing that `r=1`, and `a1=10` fill in the blanks.